OpenStudy (richyw):

Higher-dimensional linear systems (ordinary differential equations) \[X'=\left[\begin{matrix} 0 & 1 & 0 \\ -1 & 0 &0 \\0 & 0 & -1\end{matrix}\right]X\]

OpenStudy (richyw):

So clearly I have eigenvalues 1, and \(\pm i\). Now what do I do?

OpenStudy (richyw):

my book jumps straight to the solution satisfying the initial condition \(X_0=(x_0,y_0,z_0)\)\[Y(t)=x_0\left[\begin{matrix} \cos t \\ -\sin t \\0 \end{matrix}\right]+y_0\left[\begin{matrix} \sin t \\ \cos t \\0 \end{matrix}\right]+z_0\left[\begin{matrix} 0 \\ 0\\1 \end{matrix}\right]\] no explanation (and what is X(t)?)

OpenStudy (richyw):

also never purchase hirsch's textbook. It's terrible and lacking so much content that they use like thin cardstock for the pages to make it thick enough! (don't even think about trying to lay this book open while you take notes!)

OpenStudy (anonymous):

0+-i = rotation .. by de-moivre conversion of imaginary to argand

OpenStudy (richyw):

sorry what are you saying?

OpenStudy (richyw):

oops and I meant -1 for the real eignevalue

OpenStudy (richyw):

nevermind can see that I missread the last term. now I can clearly see what is going on. thanks!